# Surface topological quantum criticality: Conformal manifolds and Discrete Strong Coupling Fixed Points

### v1

@@ III.3.1 Effective field theory approach
We use the effective field theory(EFT) approach to derive the unitary operator. Here, we look at the effective low-energy model of the Hamiltonian given in Eqn.1. By expanding to first order in k around all the time-reversal (TR) invariant points in the bulk Brillouin zone (BZ), only the mass terms in the Hamiltonian persist. Since the effective model only includes TR invariant points from the bulk BZ, it becomes simpler to construct a subspace of the BZ spanned solely by the TR invariant momenta. Consequently, the effective Hamiltonian in momentum space reduces to an $8\times 8$ matrix with the diagonal entries corresponding to the respective mass gaps.
To deform the surface state, we require the mass to be flipped only on the $k_{z}=\pi$ plane as discussed in section.IIIB. So, to look for possible ways to flip the masses, we construct an effective $4\times 4$ Hamiltonian matrix in the subspace spanned by the four TRI points in the $k_{z}=\pi$ plane (see fig.3). Then, the effective $4\times 4$ Hamiltonian has a diagonal representation as follows:
@@
We define another set of Pauli matrices $\sigma_{i}(i=x,y,z)$ in this subspace. We can recast the Hamiltonian defined in Eqn.4 in terms of $\Sigma_{i}$’s and $\sigma_{i}$’s as,
@@
Here, $\sigma_{y}$ signifies a coupling between the mass points separated by the momentum $\textbf{K}=(\pi,\pi,0)$. Applying the unitary transformation to the Hamiltonian in Eqn.6, we see that the signs of the four mass gaps flip when $\lambda$ approaches one.
@@
The interchange of the mass points is evident in the above Hamiltonian matrix.
@@
Notably, for any intermediate value of $\lambda$, the lattice translation and the inversion symmetries are broken. The breaking of translation symmetry is because we introduced the coupling between the momentum states. Furthermore, since $U_{1}(\lambda)$ contains the $\gamma$ operator (see Eqn.8), which anti-commutes with the mass term $\beta$, it also breaks the inversion symmetry. However, when $\lambda=1$, these symmetries are restored. Restoring inversion symmetry is crucial since inversion symmetry is essential for expressing the topological invariants as functions of the parity eigenvalues at each time-reversal invariant (TRI) point.
III.3.2 Microscopic model
@@
We consider the Hamiltonian given in Eqn.1. We shall take an isotropic model by setting $t_{x}=t_{y}=t_{z}=t$. In the above representation, the bulk Hamiltonian takes the form,
@@
Recall that while deriving the unitary operator using the effective field theory, we reduced the full $8\times 8$ mass matrix to a simple $4\times 4$ matrix by assuming that the unitary evolution affects only the four TRI points in the $k_{z}=\pi$ plane(see Eqn.4). We implicitly assumed this while deriving the unitary operator. However, when we construct a microscopic lattice model for the unitary operator, we must introduce a momentum-dependent function in the unitary operator such that the action of the unitary operator is restricted to the top four mass points. To realize this, we introduce the following Kronecker delta function into the unitary operator,
@@
However, the resulting Hamiltonian will lead to an infinite-range hopping effect in the bulk lattice. For numerical simulations, it is beneficial to replace the Kronecker delta function in Eqn.15 with a smooth function, such as a Gaussian function as,
@@
We now insert $\phi_{\textbf{k}}$ defined in Eqn.16 into the unitary operator in Eqn.11 to restrict the action of unitary evolution to the $k_{z}=\pi$ plane of the bulk BZ. The modified unitary operator now reads,
@@
The Hamiltonian in Eqn.III.3.2 under the unitary operation $U(k_{z},\lambda)$ takes the form,
@@
Figure 4: a) Schematic of two different pathways connecting the $\mathcal{N}=1$ and $\mathcal{N}=3$ TI surfaces for the same bulk topological invariant. Along the black line, all symmetries are preserved, but the $\mathcal{N}=1$ surface can be connected to $\mathcal{N}=3$ surface only by a gap closing or via a critical point. On the other hand, along the semi-circle, the $\mathcal{N}=1$ surface can be smoothly connected to the $\mathcal{N}=3$ surface without closing the bulk gap. However, several crystal symmetries are broken. Also shown is the surface spectrum when $\theta=0$ (b) and $\theta=\pi$ (c) by solving the Hamiltonian in Eqn.18. $\textbf{Q}_{i}$’s ($i=1,2,3,4$) are the TRI points in the surface BZ. (See Fig.1 for the exact definition).
@@
When $\lambda=1$ and on taking the limit of $\epsilon\rightarrow 0$, the Hamiltonian in Eqn.18 becomes,
@@
Since $\phi_{\textbf{k}}=\delta_{k_{z},\pi}$, the mass signs at the four TRI points in the $k_{z}=\pi$ plane are flipped when $\lambda$ is tuned to one. Therefore, as evident from the fig.2, if the Hamiltonian before the unitary transformation is in a strong TI phase with a single Dirac cone, then after the unitary transformation, we arrive at the strong TI phase of the same set of topological invariants but with three Dirac cones without closing the gap.
@@
Here, we proposed a possible unitary operation that could smoothly connect the distinct surface states of TI with identical topological bulk. Theoretically, such an adiabatic path should exist because both surfaces are in the same bulk topological phase. We showed here that such a path indeed exists. However, the prize of introducing such a unitary deformation is that it lowers the symmetry of the system. The translation and the inversion symmetry are broken for intermediate values of the parameter $\lambda$ between 00 and $1$. In addition, the unitary operation introduces infinite-range hopping in the bulk lattice.
A practical realization of such a unitary path would be challenging. Therefore, we argue that the number of surface fermion flavors, $\mathcal{N}$, can be stabilized by imposing crystal symmetry constraints. This gives us the confidence to study the effect of strong interactions on the TI surface with $\mathcal{N}$ flavors of surface fermions. Unless the interactions break the symmetries mentioned above or introduce long-range hopping, the number of flavors $\mathcal{N}$ should remain topologically stable.
Although we have demonstrated this point in the case of a strong topological insulator (TI), the same principles also apply to a weak TI. Consider the Hamiltonian in Eqn.1 representing a weak TI phase with two Dirac cones ($\mathcal{N}=2$). Upon applying the unitary operation $U(\lambda)$, we observe that at $\lambda=1$, the TI surface retains two Dirac cones. Yet, they are located at complementary points in the surface Brillouin zone (BZ). For instance, if the Dirac cones were at $(0,0)$ and $(\pi,0)$ for $\lambda=0$, then after applying the unitary deformation at $\lambda=1$, the Dirac cones would be found at $(\pi,\pi)$ and $(0,\pi)$. For this reason, the location of the Dirac cones before and after the transformation are complementary.
IV Interacting dynamics of the TI surface with $\mathcal{N}$-flavors of surface Dirac fermions
@@
IV.1 Uniqueness of the gapless surface states
@@
In the $\alpha-\beta$ representation introduced in this article, $\Gamma_{s}=\tau_{y}s_{z}$, where $\tau_{y}$ is the Pauli matrix for the orbital part and $s_{z}$ is the Pauli matrix for spins (see Eqn.1).
@@
This is rather an intrinsic property of the topological insulator, disregarding the detailed structure of the surface Brillouin zone. Now let us compare it with a gapless Dirac fermion in a bulk 2D lattice. Each Dirac point will always have both positive and negative pseudo-chiralities. Therefore, a 2D lattice would always have an equal number of eigenstates for each eigenvalue of the $\Gamma_{s}$ operator. So a bulk 2D lattice cannot be associated with a specific chirality. This is an intrinsic property of lattices due to the well-known fermion-doubling theory[40, 41].
@@
IV.2 Effective 2D interacting model
In section III, we started with a 3D Dirac lattice model suitable for a cubic lattice and solved it numerically to determine the surface states. However, since we want to study the effects of interactions on the TI surface, we need to better understand the analytical form of the surface states, at least near the corresponding TRI points. As discussed at the beginning of this section, we cannot simply start with a 2D lattice model to discuss surface fermions.
@@ IV.2 Effective 2D interacting model
Whether a Dirac cone is present or absent at the given TRI point is solely determined by the TI bulk, as we saw in detail in section III. Therefore, one must start with the bulk topological phase and find out the precise value of $\varepsilon_{i},(i=1,2,3,4)$, and the locations of Dirac cones in the surface BZ to construct the effective surface theory.
@@
Here $V$ is the attractive interaction strength in the 3D bulk. It is non-zero within the energy scale $[-\Lambda,\Lambda]$, where $\Lambda$ is the UV cut-off scale of the interaction. Here $i,j$ runs over the surface fermion flavors, subject to their presence. $m_{i}$ and $m_{j}$ are the bulk mass gaps of the Dirac cones between which the scattering of the singlet pair of fermions occurs due to effective attractive interaction. In the effective theory, the singlet pair of fermions from a surface Dirac cone can scatter to either the same or to another Dirac cone (if present), the strength of which is determined by the respective bulk mass gaps assigned to each of the Dirac cones. See Fig.5 for a schematic picture of the scattering.
IV.3 Symmetries of the free surface fermions
Let us now study the symmetries of the non-interacting part of the theory defined in Eqn.LABEL:Lfrfmn. A surface with $\mathcal{N}$ Dirac cones can be pictured as a theory of $\mathcal{N}$ distinct flavors of 2-component Dirac fermions. We could define an internal flavor space of $\mathcal{N}$-dimensions. In this flavor space, we define a $\mathcal{N}$ component spinor as $\Psi^{\dagger}=\left(\begin{array}[]{cccc}\psi^{\dagger}_{1}&\psi^{\dagger}_{2}&\psi^{\dagger}_{3}&...\end{array}\right)$. Here $\psi^{\dagger}_{i}(i=1,2,..\mathcal{N})$ is a 2-component spinor (see Eqn.21). In this section, we shall show that the free fermions can be $invariant$ under the $SU(\mathcal{N})$ rotations of the Dirac spinor $\Psi$ in the flavor space if a proper spinor-base is chosen.
@@ IV.3 Symmetries of the free surface fermions
Consider a Dirac cone at the surface TRI point $\textbf{Q}_{i}$ (see Fig.5 for its definition). We look for a possible $SU(2)$ rotation in the spinor space such that the helicity of the Dirac cone at $\textbf{Q}_{i},i=1,2,3,4$ is made equal to $\textbf{Q}_{1}$. Let $U_{i}$ be the unitary matrix defined for the Dirac cone at $\textbf{Q}_{i}$ which realizes this transformation. $U_{i}$’s would have the following definitions,
@@
In general, a unitary matrx $U_{i}$, $i=2,3,4$ or $is_{\alpha}(\alpha=x,y,z)$ realizes an $SU(2)$ spinor rotation about the $\alpha$-direction by an angle $\pi$ for a surface fermion at the TRI point $\textbf{Q}_{i}$. Rotating each flavor of the surface fermions independently as $\psi_{i}\rightarrow U_{i}\psi_{i}$, the free-fermion dynamics of all the Dirac cones appear to be identical.
@@
As a result, we express the free theory of surface fermions in Eqn.21 in terms of the $\mathcal{N}$-component spinor field $\Psi(\textbf{x})$ as,
@@
Here $\hat{I}_{\mathcal{N}}$ is an identity matrix in the flavor space. $\hat{h}_{1}$ is defined in Eqn.21. Also, notice that we haven’t used the term $\varepsilon_{i}$ here. This is because the helicity invariance implies the location of the surface Dirac cones is not relevant to the theory. Only the number of surface Dirac cones $\mathcal{N}$ matters.
One can see that the free theory of surface fermions in Eqn.24 is invariant under $SU(\mathcal{N})$ flavor rotations of the Dirac spinor. The single-particle Green’s function in the flavor space has the simple form of
@@
IV.4 Symmetries of the interaction matrix $V_{ij}$

### v3

@@ IV.3.1 Effective field theory approach
We use the effective field theory(EFT) approach to derive the unitary operator. Here, we look at the effective low-energy model of the Hamiltonian given in Eqn.2. By expanding to first order in k around all the time-reversal (TR) invariant points in the bulk Brillouin zone (BZ), only the mass terms in the Hamiltonian persist. Since the effective model only includes TR invariant points from the bulk BZ, it becomes simpler to construct a subspace of the BZ spanned solely by the TR invariant momenta. Consequently, the effective Hamiltonian in momentum space reduces to an $8\times 8$ matrix with the diagonal entries corresponding to the respective mass gaps.
To deform the surface state, we require the mass to be flipped only on the $k_{z}=\pi$ plane as discussed in section.IVB. So, to look for possible ways to flip the masses, we construct an effective $4\times 4$ Hamiltonian matrix in the subspace spanned by the four TRI points in the $k_{z}=\pi$ plane (see fig.3). Then, the effective $4\times 4$ Hamiltonian has a diagonal representation as follows:
@@
We define another set of Pauli matrices $\sigma_{i}(i=x,y,z)$ in this subspace. We can recast the Hamiltonian defined in Eqn.5 in terms of $\Sigma_{i}$’s and $\sigma_{i}$’s as,
@@
Here, $\sigma_{y}$ signifies a coupling between the mass points separated by the momentum $\textbf{K}=(\pi,\pi,0)$. Applying the unitary transformation to the Hamiltonian in Eqn.7, we see that the signs of the four mass gaps flip when $\lambda$ approaches one.
@@
The interchange of the mass points is evident in the above Hamiltonian matrix. Consequently, this operation results in a change of the bulk topology. This is no surprise as the three weak topological invariants $v_{i}$($i=1,2,3$) are robust only if the crystal translation symmetry is respected. Here, we broke the translation symmetry by introducing the momentum coupling. So the three invariants can change while keeping the strong topological invariant $v_{0}$ unchanged.
@@
Notably, for any intermediate value of $\lambda$, the lattice translation and the inversion symmetries are broken. The breaking of translation symmetry is because we introduced the coupling between the momentum states. Furthermore, since $U_{1}(\lambda)$ contains the $\gamma$ operator (see Eqn.9), which anti-commutes with the mass term $\beta$, it also breaks the inversion symmetry. However, when $\lambda=1$, these symmetries are restored. Restoring inversion symmetry is crucial since inversion symmetry is essential for expressing the topological invariants as functions of the parity eigenvalues at each time-reversal invariant (TRI) point.
IV.3.2 Microscopic model
@@
We consider the Hamiltonian given in Eqn.2. We shall take an isotropic model by setting $t_{x}=t_{y}=t_{z}=t$. In the above representation, the bulk Hamiltonian takes the form,
@@
Recall that while deriving the unitary operator using the effective field theory, we reduced the full $8\times 8$ mass matrix to a simple $4\times 4$ matrix by assuming that the unitary evolution affects only the four TRI points in the $k_{z}=\pi$ plane(see Eqn.5). We implicitly assumed this while deriving the unitary operator. However, when we construct a microscopic lattice model for the unitary operator, we must introduce a momentum-dependent function in the unitary operator such that the action of the unitary operator is restricted to the top four mass points. To realize this, we introduce the following Kronecker delta function into the unitary operator,
@@
However, the resulting Hamiltonian will lead to an infinite-range hopping effect in the bulk lattice. For numerical simulations, it is beneficial to replace the Kronecker delta function in Eqn.16 with a smooth function, such as a Gaussian function as,
@@
We now insert $\phi_{\textbf{k}}$ defined in Eqn.17 into the unitary operator in Eqn.12 to restrict the action of unitary evolution to the $k_{z}=\pi$ plane of the bulk BZ. The modified unitary operator now reads,
@@
The Hamiltonian in Eqn.IV.3.2 under the unitary operation $U(k_{z},\lambda)$ takes the form,
@@
Figure 4: a) Schematic of two different pathways connecting the $\mathcal{N}=1$ and $\mathcal{N}=3$ TI surfaces for the same bulk topological invariant. Along the black line, all symmetries are preserved, but the $\mathcal{N}=1$ surface can be connected to $\mathcal{N}=3$ surface only by a gap closing or via a critical point. On the other hand, along the semi-circle, the $\mathcal{N}=1$ surface can be smoothly connected to the $\mathcal{N}=3$ surface without closing the bulk gap. However, several crystal symmetries are broken. Also shown is the surface spectrum when $\theta=0$ (b) and $\theta=\pi$ (c) by solving the Hamiltonian in Eqn.19 for the bulk topological phase $(1;1,1,1)$. $\textbf{Q}_{i}$’s ($i=1,2,3,4$) are the TRI points in the surface BZ. (See Fig.1 for the exact definition).
@@
When $\lambda=1$ and on taking the limit of $\epsilon\rightarrow 0$, the Hamiltonian in Eqn.19 becomes,
@@
Since $\phi_{\textbf{k}}=\delta_{k_{z},\pi}$, the mass signs at the four TRI points in the $k_{z}=\pi$ plane are flipped when $\lambda$ is tuned to unity. Therefore, as evident from the fig.2, if the Hamiltonian before the unitary transformation is in a strong TI phase with a single Dirac cone, then after the unitary transformation, we arrive at the strong TI phase of the same set of topological invariants but with three Dirac cones without closing the gap.
@@
Recall that the unitary operation $U_{1}(k_{z},\lambda)$ flips the mass signs in the $k_{z}=\pi$ plane; however, it also interchanges the mass points separated by momentum K (see Eqn.10). As a result, the unitary operation changes the bulk topological invariants, even though it deforms a single Dirac cone surface state to three Dirac cone surface states (see Fig.12). Details are moved to the Appendix.B. Just like $U(k_{z},\lambda)$, $U_{1}$ is also a non-local unitary operation and breaks the translation and inversion symmetries of the bulk for intermediate values of $\lambda$.
Here, we proposed two non-local unitary operations that could connect distinct surface states but without breaking the protecting symmetry. While the first one $U_{1}(k_{z},\lambda)$(Eqn.18a) changes the bulk topological invariants, the final unitary $U(k_{z},\lambda)$(Eqn.18c) connects distinct surface states without changing the topological bulk. Thus, we showed that an adiabatic path exists that could connect distinct surface states without closing the bulk gap and without breaking the protecting symmetry. However, the cost of introducing such a unitary deformation is that it lowers the symmetry of the system. The translation and the inversion symmetry are broken for intermediate values of the parameter $\lambda$ between 00 and $1$. In addition, the unitary operation is non-local in the sense that it introduces infinite-range hopping in the bulk lattice.
A practical realization of such a non-local unitary path would be challenging. Therefore, we argue that the number of surface fermion flavors, $\mathcal{N}$, can be stabilized by imposing crystal symmetry constraints. This gives us the confidence to study the effect of strong interactions on the TI surface with $\mathcal{N}$ flavors of surface fermions. Unless the interactions break the symmetries mentioned above or introduce infinite-range hopping, the number of flavors $\mathcal{N}$ should remain topologically stable.
Although we have demonstrated this point in the case of a strong topological insulator (TI), the same principles also apply to a weak TI. Consider the Hamiltonian in Eqn.2 representing a weak TI phase with two Dirac cones ($\mathcal{N}=2$). Upon applying the unitary operation $U(\lambda)$, we observe that at $\lambda=1$, the TI surface retains two Dirac cones. Yet, they are located at complementary points in the surface Brillouin zone (BZ). For instance, if the Dirac cones were at $(0,0)$ and $(\pi,0)$ for $\lambda=0$, then after applying the unitary deformation at $\lambda=1$, the Dirac cones would be found at $(\pi,\pi)$ and $(0,\pi)$. For this reason, the locations of the Dirac cones before and after the transformation are complementary.
V Interacting dynamics of the TI surface with $\mathcal{N}$-flavors of surface Dirac fermions
@@
V.1 Uniqueness of the gapless surface states
@@
In the $\alpha-\beta$ representation introduced in this article, $\Gamma_{s}=\tau_{y}s_{z}$, where $\tau_{y}$ is the Pauli matrix for the orbital part and $s_{z}$ is the Pauli matrix for spins (see Eqn.2).
@@
This is rather an intrinsic property of the topological insulator, disregarding the detailed structure of the surface Brillouin zone. Now, let us compare it with a gapless Dirac fermion in a bulk 2d lattice. Each Dirac point will always have both positive and negative pseudo-chiralities. Therefore, a 2d lattice would always have an equal number of eigenstates for each eigenvalue of the $\Gamma_{s}$ operator. So a bulk 2d lattice cannot be associated with a specific chirality. This is an intrinsic property of lattices due to the well-known fermion-doubling theory[42, 43].
@@
V.2 Effective 2d interacting model
In section IV, we started with a 3d Dirac lattice model suitable for a cubic lattice and solved it numerically to determine the surface states. However, since we want to study the effects of interactions on the TI surface, we need to better understand the analytical form of the surface states, at least near the corresponding TRI points. As discussed at the beginning of this section, we cannot simply start with a 2d lattice model to discuss surface fermions.
@@ V.2 Effective 2d interacting model
Whether a Dirac cone is present or absent at the given TRI point is solely determined by the TI bulk, as we saw in detail in section IV. Therefore, one must start with the bulk topological phase and find out the precise value of $\varepsilon_{i},(i=1,2,3,4)$, and the locations of Dirac cones in the surface BZ to construct the effective surface theory.
@@
Here, $V$ is the attractive interaction strength in the 3D bulk. It is non-zero within the energy scale $[-\Lambda,\Lambda]$, where $\Lambda$ is the UV cut-off scale of the interaction. Here $i,j$ runs over the surface fermion flavors, subject to their presence. $m_{i}$ and $m_{j}$ are the bulk mass gaps of the Dirac cones between which the scattering of the singlet pair of fermions occurs due to effective attractive interaction. In the effective theory, the singlet pair of fermions from a surface Dirac cone can scatter to either the same or to another Dirac cone (if present), the strength of which is determined by the respective bulk mass gaps assigned to each of the Dirac cones. See Fig.5 for a schematic picture of the scattering.
V.3 Symmetries of the free surface fermions
Let us now study the symmetries of the non-interacting part of the theory defined in Eqn.LABEL:Lfrfmn. A surface with $\mathcal{N}$ Dirac cones can be pictured as a theory of $\mathcal{N}$ distinct flavors of 2-component Dirac fermions. We could define an internal flavor space of $\mathcal{N}$-dimensions. In this flavor space, we define a $\mathcal{N}$ component spinor as $\Psi^{\dagger}=\left(\begin{array}[]{cccc}\psi^{\dagger}_{1}&\psi^{\dagger}_{2}&\psi^{\dagger}_{3}&...\end{array}\right)$. Here $\psi^{\dagger}_{i}(i=1,2,..\mathcal{N})$ is a 2-component spinor (see Eqn.22). In this section, we shall show that the free fermions can be $invariant$ under the $SU(\mathcal{N})$ rotations of the Dirac spinor $\Psi$ in the flavor space if a proper spinor-base is chosen.
@@ V.3 Symmetries of the free surface fermions
Consider a Dirac cone at the surface TRI point $\textbf{Q}_{i}$ (see Fig.5 for its definition). We look for a possible $SU(2)$ rotation in the spinor space such that the helicity of the Dirac cone at $\textbf{Q}_{i},i=1,2,3,4$ is made equal to $\textbf{Q}_{1}$. Let $U_{i}$ be the unitary matrix defined for the Dirac cone at $\textbf{Q}_{i}$, which realizes this transformation. $U_{i}$’s would have the following definitions,
@@
In general, a unitary matrix $U_{i}$, $i=2,3,4$ or $is_{\alpha}(\alpha=x,y,z)$ realizes an $SU(2)$ spinor rotation about the $\alpha$-direction by an angle $\pi$ for a surface fermion at the TRI point $\textbf{Q}_{i}$. Rotating each flavor of the surface fermions independently as $\psi_{i}\rightarrow U_{i}\psi_{i}$, the free-fermion dynamics of all the Dirac cones appear to be identical.
@@
As a result, we express the free theory of surface fermions in Eqn.22 in terms of the $\mathcal{N}$-component spinor field $\Psi(\textbf{x})$ as,
@@
Here $\hat{I}_{\mathcal{N}}$ is an identity matrix in the flavor space. $\hat{h}_{1}$ is defined in Eqn.22. Also, notice that we haven’t used the term $\varepsilon_{i}$ here. This is because the helicity invariance implies the location of the surface Dirac cones is not relevant to the theory. Only the number of surface Dirac cones $\mathcal{N}$ matters.
One can see that the free theory of surface fermions in Eqn.25 is invariant under $SU(\mathcal{N})$ flavor rotations of the Dirac spinor. The single-particle Green’s function in the flavor space has the simple form of
@@
V.4 Symmetries of the interaction matrix $V_{ij}$